# Modeling and Numerical Analysis of Kangaroo Lower Body based on Constrained Dynamics of Hybrid Serial-Parallel Floating-Base Systems

### 07 Dec 2023 v1

@@ Abstract
keywords:
Humanoid robots, kinematics, dynamics, serial-parallel hybrid chains
[…]
1 Introduction
@@
In the past, the design of biped humanoid robots has predominantly relied on well-established systems, such as the HRP series, kaneko2002design ; kaneko2008humanoid , or ASIMO, hirose2007honda , characterized by stiff position control, serial kinematics, electrical rotary actuators, and high reductions gearboxes, (see also tellez2008reem ; park2007mechanical ), with very few exceptions, e.g. LOLA lohmeier2009humanoid and TULIP hobbelen2008 . While these robots demonstrated the ability to perform complex motions, their interaction with the environment and agility were considerably limited.
@@
During the DARPA Robotics Challenge (DRC), humanoid torque-controlled robots with high-power capabilities were specifically designed for intervention in disaster scenarios spenko2018darpa , showcasing significant advancements, particularly in terms of interaction capabilities. However, these systems exhibited certain limitations, including slow movements and low-impact resilience  Tsagarakis:2017 ; radford2015valkyrie ; kojima2015development ; stasse2017talos . The sluggishness of these robots can be attributed to the heavy weight of the platforms and their high inertia, which stemmed from the utilization of actuators positioned near the moving joints. Furthermore, the incorporation of gearboxes with high reduction ratios, such as harmonic drives, contributed to their low-impact resilience. Another common characteristic of these systems was the use of torsional elements or strain gauges for joint torque measurement, rather than relying on current measurements. This approach was primarily necessitated by the high-friction nature of the high-reduction gearboxes, which made it challenging to estimate and compensate.
After the DRC, novel humanoid platforms have been developed following different design paradigms, in particular: relocating actuators to improve mass distribution, especially on the legs, and employing different actuation units than classical electrical actuators with harmonic drivers, to enhance impact robustness and torque/force control. Remarkable examples of this new generation of biped robots are the new Atlas from Boston Dynamics, see guizzo2019leaps , Cassie/Digit from Agility Robotics hurst2019walk , and UCLA’s Artemis zhu2023design . While Atlas relies on hydraulic actuation to achieve robustness and high-power impulsive motions, Cassie/Digit and Artemis use high-torque motors paired with high-efficiency low-reduction-ratio gearboxes located near the main body of the robot, with the motion transmitted through a series of closed linkages to reduce foot inertia.
@@
The paper is organized as follows: Section 2 presents related works on humanoid bipedal platforms using closed linkages, their analysis, and control methods. Section 3 introduces the background on kinematic and dynamic modeling of floating-base robots presenting serial-parallel hybrid linkages. Section 4 presents the Kangaroo robot and its lower-body kinematic structure, together with an analysis of the closed sub-mechanisms present in the legs, including joint limits and non-linear transmission effects. In Section 5, we carry out the comparison against the TALOS biped robot. The full-model validation and preliminary experiments with a prototype of the Kangaroo robot are reported in Section 6. Finally, Section 7 closes the paper with final remarks and future development.
2 Related Works
@@
The main reason to adopt serial mechanisms relies on their straightforward structure, augmented workspace, and simple manufacturing and maintenance processes, therefore representing most of the time the state-of-the-art in robotics systems. For these reasons, most bipedal robots’ legs consist of 6 DOFs serial mechanisms with actuators distributed along the whole kinematic chain, from the hip to the ankle, see for example kaneko2002design ; hirose2007honda ; tellez2008reem ; park2007mechanical . However, despite the aforementioned advantages, these types of robot architectures generally present only limited precision, low structural stiffness, and poor dynamic characteristics, related to mass and inertia distribution. In particular, the latter plays a fundamental role in agile locomotion and impact handling and mitigation. On the other hand, a parallel robot can provide higher stiffness, speed, accuracy, and payload capacity, at the cost of a reduced workspace and complex geometry, requiring careful analysis and control. kumar2020survey defines a series-parallel hybrid robot as a robot constituted by a serial or tree-type combination of serial and parallel mechanisms, combining the advantages of both worlds, but also inheriting their kinematic complexities.
@@
The last few decades of research in humanoid bipedal robotics have shown that achieving high dynamic performance requires a stiff structure and good mass distribution stasse2019overview . These characteristics can be achieved using series-parallel hybrid structures. In fact, more recently, several research works started to introduce mechanical leg designs that are based on series-parallel hybrid kinematics. In some cases, sub-mechanisms are based on parallel kinematics to achieve the lightweight, modular, and compact design of sub-parts. A typical case is the ankle, where motors are relocated closer to the knee by employing a four-bar linkage mechanism with a reduction of inertia at the leg end-effector, as in Tsagarakis:2017 ; radford2015valkyrie ; Ruscelli18 . A first example of a humanoid bipedal robot strongly based on series-parallel hybrid kinematics chains is LOLA, from lohmeier2009humanoid , which legs were designed mixing rotational, linear actuators, and parallel/differential transmissions. Other examples of humanoid bipedal systems where series-parallel hybrid chains were extensively used for the upper and lower body are HYDRA, Kaminaga16 , SAFFIR, lahr2016biologically , CARL, schutz2017carl , and Disney’s HYBRID LEG, gim2018design , to name a few.
@@
From the software and control point of view, the parallel mechanism is often handled by relying on an abstraction level that separates the parallel-closed chain computation into specialized functions (e.g. transmissions in ros_control chitta2017ros_control ), or approximates the system as a serial chain.
@@
In mronga2021whole , is presented a Quadratic Programming (QP)-based framework for whole-body control taking into account closed kinematic loops. The authors show how approaches that are based on serial-chain abstraction and separated into specialized functions lead to theoretical and practical disadvantages, in particular:
@@
The proposed framework is applied to a full-size humanoid robot with multiple series-parallel hybrid chains, named RH5, see Kumar20 ; ebetaer2021design .
Recently, agile and dynamic walking has been achieved with platforms based on parallel mechanism legs, e.g. ATRIAS, grimes2012design , and series-parallel hybrid mechanism, e.g. Cassie and Digit reher2019dynamic . ATRIAS leg design is based on a four-bar linkage driven by 2 motors for leg extension/retraction in common mode, and swing in differential mode. The design of ATRIAS has evolved in Cassie/Digit, employing a series-parallel hybrid mechanism modeled by cutting the loops and adding consistency constraints, see apgar2018fast .
In gim2018design is proposed an IK approach to retarget motions computed from a serial model onto a hybrid one to control the Disney’s HYBRID LEG, see also schumacher2021versatile . However, the proposed approach is only kinematics-based.
The work in kumar2020survey reports (in Table 1) a comparison among different series-parallel hybrid humanoid robots, considering the number of parallel sub-mechanism modules and the provided free DOFs w.r.t. the total number of free DOFs. The same table is reported in Table 1, considering only humanoid robots and augmented with the data for the Kangaroo robot. It is worth noticing that among all the considered humanoid bipedal systems, Kangaroo is the only one relocating the ankle’s actuators at the back of the top rear femur, hence not presenting any actuator nor electronics under the knee. This unique feature makes Kangaroo a platform particularly resilient against impacts and potentially ideal for performing agile and dynamic locomotion.
Table 1: Overview of parallel sub-mechanism based modules with different complexities in serial-parallel hybrid robots
@@
LOLA, lohmeier2009humanoid 2 2 25 6
AILA, lemburg2011aila 2 2 20 6
Valkyrie, radford2015valkyrie 5 35 10
TORO, englsberger2014overview 2 27 2
THOR, lee2014thesis 4 4 30 12
SAFFIR, lahr2014thesis 4 4 30 12
LARMBOT, cafolla2016larmbot 2 1 22 10
TALOS, stasse2017talos 2 27 2
RH5, ebetaer2021design 5 5 32 15
Disney Biped, gim2018design 2 12 12
@@
3 Background on Modeling Serial-Parallel Hybrid Chains in Floating-Base Systems
Classical approaches in modeling parallel linkages are based on geometrical analysis, leveraging case-specific analytical expressions for the linkage kinematics, as in kumar2019kinematic where the closed-form solution of a 2SPRR-U humanoid ankle is presented. This solution permits embedding the closed loop constraints directly in the equation of motions eventually saving computational cost and improving numerical accuracy. However, such closed-form solutions are not always available and do not generalize to arbitrary serial-parallel hybrid mechanisms.
According to laulusa2008review , a more comprehensive and promising method entails modeling these linkages as constrained multi-body systems. This approach allows for the consideration of intricate geometries and dynamic effects associated with multiple serial-parallel hybrid chains. While the direct map between the linkage parameters and the resulting motion is no longer expressed analytically, it can still be determined through numerical computation. In this section, we recall this approach, including floating-base systems incorporating generic series-parallel hybrid chains, with the aim of using it for the modeling and analysis of the Kangaroo robot.
@@ 3 Background on Modeling Serial-Parallel Hybrid Chains in Floating-Base Systems
3.1 Modeling of Serial-Parallel Hybrid Chains
@@
Figure 2: 1DOF RRPR closed mechanism. It can be modeled using 2 linear constraints.
@@
with $\bm{\theta}_{u}\in\mathbb{R}^{m}$ the passive DOFs and $\bm{\theta}_{a}\in\mathbb{R}^{n-m}$ the actuated DOF, $m<n$. For the 1DOF RRPR mechanism considered, we have in particular $n=3$ and $m=2$.
@@
where ${{}^{a}\mathbf{v}_{a,u}}$ is the velocity along the constrained directions of the frame $\mathcal{F}_{u}$ w.r.t. the frame $\mathcal{F}_{a}$ expressed in frame $\mathcal{F}_{a}$ (please refer to Appendix A. Computation of Relative Kinematics on how to compute relative kinematics quantities). It is possible to divide the constraint Jacobian into its actuated and passive parts:
@@
$\bm{\dot{\theta}}_{u}$ $=-\mathbf{J}_{l,u}\left(\bm{\theta}\right)^{-1}\mathbf{J}_{l,a}\left(\bm{\theta}\right)\bm{\dot{\theta}}_{a}=\mathbf{J}_{m}\left(\bm{\theta}\right)\bm{\dot{\theta}}_{a},$ (3.1)
@@
please refer to Appendix A2. Relative Acceleration and Acceleration Bias for the computation of the relative bias term $\mathbf{\dot{J}}_{l}(\bm{\theta},\bm{\dot{\theta}})\bm{\dot{\theta}}$. Splitting (8) into its actuated and passive parts leads to:
@@
$=\mathbf{J}_{m}\left(\bm{\theta}\right)\bm{\ddot{\theta}}_{a}-\mathbf{J}_{l,u}\left(\bm{\theta}\right)^{-1}\mathbf{\dot{J}}_{l}(\bm{\theta},\bm{\dot{\theta}})\bm{\dot{\theta}}.$ (3.1)
The equations of motions consist of a constrained dynamic, see carpentier_hal-03271811 , and it can be written as:
@@
$=-{\mathbf{J}_{l,a}(\bm{\theta})^{T}}\mathbf{J}_{l,u}(\bm{\theta})^{-T},$ (3.1)
where equation (3.1) permits to map generic torques on the passive DOFs into the actuated ones. The inverse dynamics can finally be computed from (12b), plugging (13):
@@
$\quad\quad-\mathbf{J}_{l,a}\left(\bm{\theta}\right)^{T}\mathbf{J}_{l,u}\left(\bm{\theta}\right)^{-T}\left(\mathbf{M}_{u}(\bm{\theta})\bm{\ddot{\theta}}+\mathbf{h}_{u}(\bm{\theta},\bm{\dot{\theta}})\right)=$ (3.1)
@@
where we have used the (3.1).
@@
Figure 3: 2DOFs U-2RRPU differential mechanism. Each parallel branch introduces 3 linear and 1 rotational constraint.
@@
$\mathbf{q}=\mathbf{P}\bm{\theta}_{u},\quad\quad\bm{\tau}_{j}=\mathbf{P}\bm{\tau}_{u},$ (21)
it is possible to compute the DIK problem w.r.t. (3.1):
@@
and the inverse torque mapping problem w.r.t. (3.1):
@@
3.2 Floating-Base Dynamics including Serial-Parallel Hybrid Chains
@@
Taking inspiration from mansard2012dedicated , a possible inverse dynamics (ID) scheme can be composed of the following three steps:
@@
The first step can be addressed using a QP, constraining the contact forces inside linearized friction cones [^3] Here we do not report further constraints which may include torque limits (by including equations (28) and (29)), joint limits, DelPrete18 , and self-collision avoidance, Khazoom22 , to name a few:
@@
Notice that the QP in (3.2) can be rewritten projecting the passive part into the actuated one and therefore considering only the actuated quantities since the passive ones are dependent variables (for more details see Appendix B. Projected Floating-Base Dynamics).
4 Analysis of Kangaroo Lower Body

### 22 Feb 2024 v2

@@ Abstract
I Introduction
@@
In the past, the design of biped humanoid robots has predominantly relied on well-established systems, such as the HRP series, [1, 2], or ASIMO, [3], characterized by stiff position control, serial kinematics, electrical rotary actuators, and high reductions gearboxes, (see also [4, 5]), with very few exceptions, e.g. LOLA [6] and TULIP [7]. While these robots demonstrated the ability to perform complex motions, their interaction with the environment and agility were considerably limited.
@@
During the DARPA Robotics Challenge (DRC), humanoid torque-controlled robots with high-power capabilities were specifically designed for intervention in disaster scenarios [8], showcasing significant advancements, particularly in terms of interaction capabilities. However, these systems exhibited certain limitations, including slow movements and low-impact resilience  [9, 10, 11, 12]. The sluggishness of these robots can be attributed to the heavy weight of the platforms and their high inertia, which stemmed from the utilization of actuators positioned near the moving joints. Furthermore, the incorporation of gearboxes with high reduction ratios, such as harmonic drives, contributed to their low-impact resilience. Another common characteristic of these systems was the use of torsional elements or strain gauges for joint torque measurement, rather than relying on current measurements. This approach was primarily necessitated by the high-friction nature of the high-reduction gearboxes, which made it challenging to estimate and compensate.
After the DRC, novel humanoid platforms have been developed following different design paradigms, in particular: relocating actuators to improve mass distribution, especially on the legs, and employing different actuation units than classical electrical actuators with harmonic drivers, to enhance impact robustness and torque/force control. Remarkable examples of this new generation of biped robots are the new Atlas from Boston Dynamics, see [13], Cassie/Digit from Agility Robotics [14], and UCLA’s Artemis [15]. While Atlas relies on hydraulic actuation to achieve robustness and high-power impulsive motions, Cassie/Digit and Artemis use high-torque motors paired with high-efficiency low-reduction-ratio gearboxes located near the main body of the robot, with the motion transmitted through a series of closed linkages to reduce foot inertia.
@@
The paper is organized as follows: Section II presents related works on humanoid bipedal platforms using closed linkages, their analysis, and control methods. Section III introduces the background on kinematic and dynamic modeling of floating-base robots presenting serial-parallel hybrid linkages. Section IV presents the Kangaroo robot and its lower-body kinematic structure, together with an analysis of the closed sub-mechanisms present in the legs, including joint limits and non-linear transmission effects. In Section V, we carry out the comparison against the TALOS biped robot. The full-model validation and preliminary experiments with a prototype of the Kangaroo robot are reported in Section VI. Finally, Section VII closes the paper with final remarks and future development.
II Related Works
@@
The main reason to adopt serial mechanisms relies on their straightforward structure, augmented workspace, and simple manufacturing and maintenance processes, therefore representing most of the time the state-of-the-art in robotics systems. For these reasons, most bipedal robots’ legs consist of 6 DOFs serial mechanisms with actuators distributed along the whole kinematic chain, from the hip to the ankle, see for example [1, 3, 4, 5]. However, despite the aforementioned advantages, these types of robot architectures generally present only limited precision, low structural stiffness, and poor dynamic characteristics, related to mass and inertia distribution. In particular, the latter plays a fundamental role in agile locomotion and impact handling and mitigation. On the other hand, a parallel robot can provide higher stiffness, speed, accuracy, and payload capacity, at the cost of a reduced workspace and complex geometry, requiring careful analysis and control. [16] defines a series-parallel hybrid robot as a robot constituted by a serial or tree-type combination of serial and parallel mechanisms, combining the advantages of both worlds, but also inheriting their kinematic complexities.
@@
The last few decades of research in humanoid bipedal robotics have shown that achieving high dynamic performance requires a stiff structure and good mass distribution [17]. These characteristics can be achieved using series-parallel hybrid structures. In fact, more recently, several research works started to introduce mechanical leg designs that are based on series-parallel hybrid kinematics. In some cases, sub-mechanisms are based on parallel kinematics to achieve the lightweight, modular, and compact design of sub-parts. A typical case is the ankle, where motors are relocated closer to the knee by employing a four-bar linkage mechanism with a reduction of inertia at the leg end-effector, as in [9, 10, 18]. A first example of a humanoid bipedal robot strongly based on series-parallel hybrid kinematics chains is LOLA, from [6], which legs were designed mixing rotational, linear actuators, and parallel/differential transmissions. Other examples of humanoid bipedal systems where series-parallel hybrid chains were extensively used for the upper and lower body are HYDRA, [19], SAFFIR, [20], CARL, [21], and Disney’s HYBRID LEG, [22], to name a few.
@@
From the software and control point of view, the parallel mechanism is often handled by relying on an abstraction level that separates the parallel-closed chain computation into specialized functions (e.g. transmissions in ros_control [23]), or approximates the system as a serial chain.
@@
In [24], is presented a Quadratic Programming (QP)-based framework for whole-body control taking into account closed kinematic loops. The authors show how approaches that are based on serial-chain abstraction and separated into specialized functions lead to theoretical and practical disadvantages, in particular:
@@
The proposed framework is applied to a full-size humanoid robot with multiple series-parallel hybrid chains, named RH5, see [25, 26].
Recently, agile and dynamic walking has been achieved with platforms based on parallel mechanism legs, e.g. ATRIAS, [27], and series-parallel hybrid mechanism, e.g. Cassie and Digit [28]. ATRIAS leg design is based on a four-bar linkage driven by 2 motors for leg extension/retraction in common mode, and swing in differential mode. The design of ATRIAS has evolved in Cassie/Digit, employing a series-parallel hybrid mechanism modeled by cutting the loops and adding consistency constraints, see [29].
In [22] is proposed an IK approach to retarget motions computed from a serial model onto a hybrid one to control the Disney’s HYBRID LEG, see also [30]. However, the proposed approach is only kinematics-based.
The work in [16] reports (in Table 1) a comparison among different series-parallel hybrid humanoid robots, considering the number of parallel sub-mechanism modules and the provided free DOFs w.r.t. the total number of free DOFs. The same table is reported in Table I, considering only humanoid robots and augmented with the data for the Kangaroo robot. It is worth noticing that among all the considered humanoid bipedal systems, Kangaroo is the only one relocating the ankle’s actuators at the back of the top rear femur, hence not presenting any actuator nor electronics under the knee. This unique feature makes Kangaroo a platform particularly resilient against impacts and potentially ideal to perform agile and dynamic locomotion.
TABLE I: Overview of parallel sub-mechanism based modules with different complexities in serial-parallel hybrid robots
@@
LOLA, [6] 2 2 25 6
AILA, [31] 2 2 20 6
Valkyrie, [10] 5 35 10
TORO, [32] 2 27 2
THOR, [33] 4 4 30 12
SAFFIR, [34] 4 4 30 12
LARMBOT, [35] 2 1 22 10
TALOS, [12] 2 27 2
RH5, [26] 5 5 32 15
Disney Biped, [22] 2 12 12
@@
III Background on Modeling Serial-Parallel Hybrid Chains in Floating-Base Systems
Classical approaches in modeling parallel linkages are based on geometrical analysis, leveraging case-specific analytical expressions for the linkage kinematics, as in [36] where the closed-form solution of a 2SPRR-U humanoid ankle is presented. This solution permits embedding the closed loop constraints directly in the equation of motions eventually saving computational cost and improving numerical accuracy. However, such closed-form solutions are not always available and do not generalize to arbitrary serial-parallel hybrid mechanisms.
According to [37], a more comprehensive and promising method entails modeling these linkages as constrained multi-body systems. This approach allows for the consideration of intricate geometries and dynamic effects associated with multiple serial-parallel hybrid chains. While the direct map between the linkage parameters and the resulting motion is no longer expressed analytically, it can still be determined through numerical computation. In this section, we recall this approach, including floating-base systems incorporating generic series-parallel hybrid chains, with the aim of using it for the modeling and analysis of the Kangaroo robot.
@@ III Background on Modeling Serial-Parallel Hybrid Chains in Floating-Base Systems
III-A Modeling of Serial-Parallel Hybrid Chains
@@
Figure 2: 1DOF RRPR closed mechanism.
@@
with $\bm{\theta}_{u}\in\mathbb{R}^{m}$ the passive DOFs and $\bm{\theta}_{a}\in\mathbb{R}^{n-m}$ the actuated DOF, $m<n$. For the 1DOF RRPR mechanism considered, we have in particular $n=3$ and with $m=2$.
@@
where ${{}^{a}\mathbf{v}_{a,u}}$ is the velocity along the constrained directions of the frame $\mathcal{F}_{u}$ w.r.t. the frame $\mathcal{F}_{a}$ expressed in frame $\mathcal{F}_{a}$ (please refer to Appendix A on how to compute relative kinematics quantities). It is possible to divide the constraint Jacobian into its actuated and passive parts:
@@
$\bm{\dot{\theta}}_{u}$ $=-\mathbf{J}_{l,u}\left(\bm{\theta}\right)^{-1}\mathbf{J}_{l,a}\left(\bm{\theta}\right)\bm{\dot{\theta}}_{a}=\mathbf{J}_{m}\left(\bm{\theta}\right)\bm{\dot{\theta}}_{a},$ (III-A)
@@
please refer to Appendix A-B for the computation of the relative bias term $\mathbf{\dot{J}}_{l}(\bm{\theta},\bm{\dot{\theta}})\bm{\dot{\theta}}$. Splitting (8) into its actuated and passive parts leads to:
@@
$=\mathbf{J}_{m}\left(\bm{\theta}\right)\bm{\ddot{\theta}}_{a}-\mathbf{J}_{l,u}\left(\bm{\theta}\right)^{-1}\mathbf{\dot{J}}_{l}(\bm{\theta},\bm{\dot{\theta}})\bm{\dot{\theta}}.$ (III-A)
The equations of motions consist of a constrained dynamic, see [38], and it can be written as:
@@
$=-{\mathbf{J}_{l,a}(\bm{\theta})^{T}}\mathbf{J}_{l,u}(\bm{\theta})^{-T},$ (III-A)
where equation (III-A) permits to map generic torques on the passive DOFs into the actuated ones. The inverse dynamics can finally be computed from (12b), plugging (13):
@@
$\quad\quad-\mathbf{J}_{l,a}\left(\bm{\theta}\right)^{T}\mathbf{J}_{l,u}\left(\bm{\theta}\right)^{-T}\left(\mathbf{M}_{u}(\bm{\theta})\bm{\ddot{\theta}}+\mathbf{h}_{u}(\bm{\theta},\bm{\dot{\theta}})\right)=$ (III-A)
@@
where we have used the (III-A).
@@
Figure 3: 2DOFs U-2RRPU differential mechanism.
@@
$\mathbf{\dot{q}}=\mathbf{P}\bm{\dot{\theta}}_{u},\quad\quad\bm{\tau}_{j}=\mathbf{P}\bm{\tau}_{u},$ (21)
it is possible to compute the DIK problem w.r.t. (III-A):
@@
and the inverse torque mapping problem w.r.t. (III-A):
@@
III-B Floating-Base Dynamics including Serial-Parallel Hybrid Chains
@@
Taking inspiration from [39], a possible inverse dynamics (ID) scheme can be composed of the following three steps:
@@
The first step can be addressed using a QP, constraining the contact forces inside linearized friction cones [^3] Here we do not report further constraints which may include torque limits (by including equations (28) and (29)), joint limits, [40], and self-collision avoidance, [41], to name a few:
@@
Notice that the QP in (III-B) can be rewritten projecting the passive part into the actuated one and therefore considering only in the actuated quantities since the passive ones are dependent variables (for more details see Appendix B).
IV Analysis of Kangaroo Lower Body